# Exact Thresholds for Noisy Non-Adaptive Group Testing

### 10 Jan 2024 v1

@@ Appendix F $-\ell$ Achievability Analysis for Near-Constant Weight Designs
Note that $D(Ce^{-\nu}\|e^{-\nu})$, as a function of $C\in(0,e^{-\nu}]$, is uniformly bounded from above. Hence, we can choose $C_{*}$ large enough so that the bound on $\tilde{P}_{1}$ in (243) dominates $2\exp(-\frac{C_{*}\ell n}{k})$ in the right-hand side of (241), and thus obtain
$\mathbb{P}\Big{(}M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}\leq\exp\Big{(}-\frac{\ell\nu n}{k}\big{[}D(Ce^{-\nu}\|e^{-\nu})+o(1)\big{]}\Big{)}.$ (244)
@@
Bounding $\mathbb{P}\big{(}\text{Bin}(\frac{Cn\nu e^{-\nu}\ell}{k},\rho)=\frac{\zeta\cdot Cn\nu e^{-\nu}\ell}{k}\big{)}$ and combining: We apply the Chernoff bound (see (23) in Lemma 1) to obtain
$\mathbb{P}\left(\text{Bin}\Big{(}\frac{Cn\nu e^{-\nu}\ell}{k},\rho\Big{)}=\frac{\zeta\cdot Cn\nu e^{-\nu}\ell}{k}\right)\leq\exp\Big{(}-\frac{Cn\nu e^{-\nu}\ell}{k}\cdot D\big{(}\zeta\|\rho\big{)}\Big{)}.$ (245)
Substituting (244) and (245) into (239) yields
$\mathbbm{E}k_{\ell,C,\zeta}\leq\binom{k}{\ell}\exp\Big{(}-\frac{\ell\nu ne^{-\nu}}{k}\underbrace{\Big{[}e^{\nu}D(Ce^{-\nu}\|e^{-\nu})+C\cdot D(\zeta\|\rho)+o(1)\Big{]}}_{:=f_{1}(C,\zeta,\rho,\nu)}\Big{)}.$ (246)
@@
Note that $f_{1}(C,\zeta,\rho,\nu)$ here has appeared in the proof of converse bound; see (192).
@@
We separately deal with the two cases $1\leq\ell\leq\log k$ and $\log k<\ell\leq\frac{k}{\log k}$ using Markov’s inequality. Specifically, following exactly the same arguments as the Bernoulli design (see (97)–(104)), we conclude the following for a given $(\ell,C,\zeta)$:
$\frac{n\nu e^{-\nu}\ell}{k}\big{[}f_{1}(C,\zeta,\rho,\nu)+o(1)\big{]}\geq(1+\eta_{1})\ell\log\frac{k}{\ell},\leavevmode\nobreak\ \text{for some }\eta_{1}>0$ (247)
@@
$\Longrightarrow$ $\leavevmode\nobreak\ \mathbb{P}\big{(}k_{\ell,C,\zeta}\geq 1\big{)}\leq\hat{P}_{\ell},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{where }\hat{P}_{\ell}=\begin{cases}(\ell\log k)^{-5},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }1\leq\ell\leq\log k\\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ k^{-5},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }\log k<\ell\leq\frac{k}{\log k}\end{cases}.$ (248)
@@
Therefore, if (247) holds, then (C1) holds for some given $(\ell,C,\zeta)$ with probability at most $\hat{P}_{\ell}$.
Condition for (C2) under given $(\ell,C,\zeta)$
As in (180), with respect to the randomness of $\mathbf{T}_{s}$, the event
$\mathscr{A}_{1}=\big{\{}N_{0}=(1+o(1))e^{-\nu}n\big{\}}$ (249)
@@
holds with $1-o(1)$ probability. In the following, for a generic event $\mathscr{E}$, an upper bound on $\mathbb{P}(\mathscr{E})$ in this part should be understood as a bound on $\mathbb{P}(\mathscr{E}\cap\mathscr{A}_{1})$, but we will only make this explicit in the concluding stage; see (293) below.
@@
We consider a fixed $(\ell,C,\zeta)$ and seek to establish the condition for (C2), i.e., (237) holding for some $\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}$ and some $\mathcal{J}^{\prime}\subset[p]\setminus s$ with $|\mathcal{J}^{\prime}|=\ell$. Note that $G_{\mathcal{J},\mathcal{J}^{\prime},1}$ in (237) is the number of positive tests in $\mathcal{N}_{0,\mathcal{J}^{\prime}}\cup\mathcal{M}_{\mathcal{J},\mathcal{J}^{\prime}}$, and thus we can further decompose $G_{\mathcal{J},\mathcal{J}^{\prime},1}$ into
$G_{\mathcal{J},\mathcal{J}^{\prime},1}=\widetilde{G}_{\mathcal{J}^{\prime},1}+U_{\mathcal{J},\mathcal{J}^{\prime},1},$ (250)
@@
where $\widetilde{G}_{\mathcal{J}^{\prime},1}$ is the number of positive tests in $\mathcal{N}_{0,\mathcal{J}^{\prime}}$ (with no dependence on $\mathcal{J}$, as reflected by the notation), and $U_{\mathcal{J},\mathcal{J}^{\prime},1}$ is the number of positive tests in $\mathcal{M}_{\mathcal{J},\mathcal{J}^{\prime}}$. Substituting (250) into (237), we can restate (C2) as follows:
$\widetilde{G}_{\mathcal{J}^{\prime},1}\geq\Big{(}\frac{1}{2}-\zeta\Big{)}\frac{Cn\nu e^{-\nu}\ell}{k}+\frac{1}{2}N_{0,\mathcal{J}^{\prime}}+\Big{(}\frac{1}{2}M_{\mathcal{J},\mathcal{J}^{\prime}}-U_{\mathcal{J},\mathcal{J}^{\prime},1}\Big{)}$ (251)
@@
$\max_{\begin{subarray}{c}\mathcal{J}^{\prime}\subset[p]\setminus s\\ |\mathcal{J}^{\prime}|=\ell\end{subarray}}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}-\frac{1}{2}N_{0,\mathcal{J}^{\prime}}\Big{)}\geq\Big{(}\frac{1}{2}-\zeta\Big{)}\frac{Cn\nu e^{-\nu}\ell}{k}-\max_{\begin{subarray}{c}\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}\\ \mathcal{J}^{\prime}\subset[p]\setminus s,|\mathcal{J}^{\prime}|=\ell\end{subarray}}M_{\mathcal{J},\mathcal{J}^{\prime}},$ (256)
@@
$\max_{\mathcal{J},\mathcal{J}^{\prime}}M_{\mathcal{J},\mathcal{J}^{\prime}}\leq\max_{\mathcal{J},\mathcal{J}^{\prime}}\widetilde{M}_{\mathcal{J},\mathcal{J}^{\prime}},$ (257)
where $(\mathcal{J},\mathcal{J}^{\prime})$ are implicitly subject to the same constraints as the last term in (256), and similarly in the developments below. We consider fixed $\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}$ and $\mathcal{J}^{\prime}\subset[p]\setminus s$ with $|\mathcal{J}^{\prime}|=\ell$. Since $\mathcal{M}_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k}$ and the test placements are uniform in $[n]$, a placement from items in $\mathcal{J}^{\prime}$ falls in $\mathcal{M}_{\mathcal{J}}$ with probability $\frac{C\nu e^{-\nu}\ell}{k}$, thus the randomness of $\mathbf{T}_{\mathcal{J}^{\prime}}$ gives $\widetilde{M}_{\mathcal{J},\mathcal{J}^{\prime}}\sim\text{Bin}(\frac{\ell\nu n}{k},\frac{C\nu e^{-\nu}\ell}{k})$. Then we can proceed analogously to the calculations in (111)–(115). In particular, our assumption $C\leq e^{\nu}$ is equivalent to $Ce^{-\nu}\leq 1$, and thus for given $(\mathcal{J},\mathcal{J}^{\prime})$ we have
$\mathbb{P}\Big{(}\widetilde{M}_{\mathcal{J},\mathcal{J}^{\prime}}\geq\frac{\ell n}{k(\log\frac{k}{\ell})^{1/2}}\Big{)}$ $\leq\mathbb{P}\Big{(}\text{Bin}\big{(}\frac{\nu\ell n}{k},\frac{\nu\ell}{k}\big{)}\geq\frac{\ell n}{k(\log\frac{k}{\ell})^{1/2}}\Big{)}$ (258)
$\leq\exp\Big{(}-\Omega\Big{(}\sqrt{\log\frac{k}{\ell}}\Big{)}\cdot\ell\log k\Big{)},$ (259)
where (259) follows from the same argument as the Bernoulli design (see (114)), with only the constant factors changing. Moreover, by repeating the union bound over $(\mathcal{J},\mathcal{J}^{\prime})$ as in (115), we obtain
$\mathbb{P}\Big{(}\max_{\mathcal{J},\mathcal{J}^{\prime}}\widetilde{M}_{\mathcal{J},\mathcal{J}^{\prime}}\geq\frac{\ell n}{k(\log\frac{k}{\ell})^{1/2}}\Big{)}\leq k^{-10\ell}$ (260)
@@
$\max_{\mathcal{J},\mathcal{J}^{\prime}}M_{\mathcal{J},\mathcal{J}^{\prime}}\leq\frac{\ell n}{k(\log\frac{k}{\ell})^{1/2}}=o\Big{(}\frac{\ell n}{k}\Big{)}.$ (261)
On this high-probability event, the necessary condition for (C2) given in (256) can be written as
$\max_{\begin{subarray}{c}\mathcal{J}^{\prime}\subset[p]\setminus s\\ |\mathcal{J}^{\prime}|=\ell\end{subarray}}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}-\frac{1}{2}N_{0,\mathcal{J}^{\prime}}\Big{)}\geq\Big{(}\frac{1}{2}-\zeta-o(1)\Big{)}\frac{Cn\nu e^{-\nu}\ell}{k}.$ (264)
Thus, to bound the probability of (C2), it suffices to bound the probability of its necessary condition (264). We proceed to study (264) by first studying a fixed $\mathcal{J}^{\prime}$ and then applying a union bound to account for the maximum.
@@
$\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}-\frac{1}{2}N_{0,\mathcal{J}^{\prime}}\geq\big{(}\frac{1}{2}-\zeta-o(1)\big{)}\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}\leq\mathbb{P}\Big{(}N_{0,\mathcal{J}^{\prime}}\leq\frac{d_{0}\nu e^{-\nu}\ell n}{k}\Big{)}$ (266)
@@
$+\sum_{\begin{subarray}{c}d_{0}<d\leq e^{\nu}\\ \frac{d\nu e^{-\nu}\ell n}{k}\in\mathbb{Z}\end{subarray}}\mathbb{P}\Big{(}N_{0,\mathcal{J^{\prime}}}=\frac{d\nu e^{-\nu}\ell n}{k}\Big{)}\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}\geq\frac{d+C(1-2\zeta)}{2}\frac{\nu e^{-\nu}\ell n}{k}\Big{|}N_{0,\mathcal{J}^{\prime}}=\frac{d\nu e^{-\nu}\ell n}{k}\Big{)}.$ (269)
@@
Case 1: $d_{0}\in[0,1]$ (i.e., $0\leq C(2\zeta-1)\leq 1-2\rho$). We will separately bound the probability terms in the right hand side of (266) and in (269).
@@
Bounding the term in (266): We first bound the term in the right-hand side of (266). By definition, ${N}_{0,\mathcal{J}^{\prime}}$ is the number of tests in $\mathcal{N}_{0}$ that contain some item from $\mathcal{J}^{\prime}$, thus on the event of $N_{0}=(1+o(1))ne^{-\nu}$ (as explained in the text below (249)), Lemma 8 gives
$\mathbb{P}\Big{(}N_{0,\mathcal{J}^{\prime}}=\frac{d\nu e^{-\nu}\ell n}{k}\Big{)}$ $\leq\exp\Big{(}-\frac{\ell\nu n}{k}\big{[}D\big{(}de^{-\nu}\|e^{-\nu}+o(1)\big{)}+o(1)\big{]}\Big{)}$ (270)
$\leq\exp\Big{(}-\frac{\ell\nu n}{k}\big{[}D\big{(}d_{0}e^{-\nu}\|e^{-\nu}\big{)}+o(1)\big{]}\Big{)}$ (271)
@@
for any $d\leq d_{0}$ such that $\frac{d\nu e^{-\nu}\ell n}{k}\in\mathbb{Z}$, where (271) follows from $d_{0}\leq 1$ and the monotonicity of $D(de^{-\nu}\|e^{-\nu})$ with respect to $d$. Therefore, the term in the right-hand side of (266) can be bounded as
$\mathbb{P}\Big{(}N_{0,\mathcal{J}^{\prime}}\leq\frac{d_{0}\nu e^{-\nu}\ell n}{k}\Big{)}$ $=\sum_{0\leq d\leq d_{0}:\frac{d\nu e^{-\nu}\ell n}{k}\in\mathbb{Z}}\mathbb{P}\Big{(}N_{0,\mathcal{J}^{\prime}}=\frac{d\nu e^{-\nu}\ell n}{k}\Big{)}$ (272)
$\leq O\big{(}\frac{\ell n}{k}\big{)}\exp\Big{(}-\frac{\ell\nu n}{k}\big{[}D\big{(}d_{0}e^{-\nu}\|e^{-\nu}\big{)}+o(1)\big{]}\Big{)}$ (273)
$\leq\exp\Big{(}-\frac{\ell\nu n}{k}\big{[}D\big{(}d_{0}e^{-\nu}\|e^{-\nu}\big{)}+o(1)\big{]}\Big{)},$ (274)
@@
where (272) follows from $d_{0}=O(1)$ (see (265)) and hence there are at most $O(\frac{\ell n}{k})$ summands, in (274) the leading factor $O(\frac{\ell n}{k})$ is absorbed into the $o(1)$ term in the exponent.
Bounding the term in (269): Analogously to (270), on the event of $N_{0}=(1+o(1))ne^{-\nu}$, Lemma 8 gives
$\mathbb{P}\Big{(}N_{0,\mathcal{J}^{\prime}}=\frac{d\nu e^{-\nu}\ell n}{k}\Big{)}\leq\exp\Big{(}-\frac{\ell\nu n}{k}\cdot\big{[}D\big{(}de^{-\nu}\|e^{-\nu}\big{)}+o(1)\big{]}\Big{)}.$ (275)
@@
Then, conditioned on $N_{0,\mathcal{J}^{\prime}}=\frac{d\nu e^{-\nu}\ell n}{k}$ we have $\widetilde{G}_{\mathcal{J}^{\prime},1}\sim\text{Bin}(\frac{d\nu e^{-\nu}\ell n}{k},\rho)$ via the randomness of $\mathbf{Z}$, and thus the condition $d>d_{0}$ (which implies $\frac{d+C(1-2\zeta)}{2}>d\rho$) allows us to apply the Chernoff bound (see (23) in Lemma 1) to obtain
$\mathbb{P}\Big{(}\text{Bin}\Big{(}\frac{d\nu e^{-\nu}\ell n}{k},\rho\Big{)}\geq\frac{d+C(1-2\zeta)}{2}\frac{\nu e^{-\nu}\ell n}{k}\Big{)}\leq\exp\left(-\frac{d\nu e^{-\nu}\ell n}{k}\cdot D\Big{(}\frac{1}{2}+\frac{C(1-2\zeta)}{2d}\big{\|}\rho\Big{)}\right).$ (276)

### 23 Oct 2024 v2

@@ Appendix F $-\ell$ Achievability Analysis for Near-Constant Weight Designs
Note that $D(Ce^{-\nu}\|e^{-\nu})$, as a function of $C\in(0,e^{-\nu}]$, is uniformly bounded from above. Hence, we can choose $C_{*}$ large enough so that the bound on $\tilde{P}_{1}$ in (F.9) dominates $2\exp(-\frac{C_{*}\ell n}{k})$ in the right-hand side of (F.7), and thus obtain
$\mathbb{P}\Big{(}M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}\leq\exp\Big{(}-\frac{\ell\nu n}{k}\big{[}D(Ce^{-\nu}\|e^{-\nu})+o(1)\big{]}\Big{)}.$ (F.10)
@@
Bounding $\mathbb{P}\big{(}\text{Bin}(\frac{Cn\nu e^{-\nu}\ell}{k},\rho)=\frac{\zeta\cdot Cn\nu e^{-\nu}\ell}{k}\big{)}$ and combining: We apply the Chernoff bound (see (A.1) in Lemma A.1) to obtain
$\mathbb{P}\left(\text{Bin}\Big{(}\frac{Cn\nu e^{-\nu}\ell}{k},\rho\Big{)}=\frac{\zeta\cdot Cn\nu e^{-\nu}\ell}{k}\right)\leq\exp\Big{(}-\frac{Cn\nu e^{-\nu}\ell}{k}\cdot D\big{(}\zeta\|\rho\big{)}\Big{)}.$ (F.11)
Substituting (F.10) and (F.11) into (F.5) yields
$\mathbbm{E}k_{\ell,C,\zeta}\leq\binom{k}{\ell}\exp\Big{(}-\frac{\ell\nu ne^{-\nu}}{k}\underbrace{\Big{[}e^{\nu}D(Ce^{-\nu}\|e^{-\nu})+C\cdot D(\zeta\|\rho)+o(1)\Big{]}}_{:=f_{1}(C,\zeta,\rho,\nu)}\Big{)}.$ (F.12)
@@
Note that $f_{1}(C,\zeta,\rho,\nu)$ here has appeared in the proof of converse bound; see (E.17).
@@
We separately deal with the two cases $1\leq\ell\leq\log k$ and $\log k<\ell\leq\frac{k}{\log k}$ using Markov’s inequality. Specifically, following exactly the same arguments as the Bernoulli design (see (D.14)–(D.21)), we conclude the following for a given $(\ell,C,\zeta)$:
$\frac{n\nu e^{-\nu}\ell}{k}\big{[}f_{1}(C,\zeta,\rho,\nu)+o(1)\big{]}\geq(1+\eta_{1})\ell\log\frac{k}{\ell},\leavevmode\nobreak\ \text{for some }\eta_{1}>0$ (F.13)
@@
$\Longrightarrow$ $\leavevmode\nobreak\ \mathbb{P}\big{(}k_{\ell,C,\zeta}\geq 1\big{)}\leq\hat{P}_{\ell},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{where }\hat{P}_{\ell}=\begin{cases}(\ell\log k)^{-5},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }1\leq\ell\leq\log k\\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ k^{-5},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }\log k<\ell\leq\frac{k}{\log k}\end{cases}.$ (F.14)
@@
Therefore, if (F.13) holds, then (C1) holds for some given $(\ell,C,\zeta)$ with probability at most $\hat{P}_{\ell}$.
Condition for (C2) for a given $(\ell,C,\zeta)$
As in (E.5), with respect to the randomness of $\mathbf{T}_{s}$, the event
$\mathscr{A}_{1}=\big{\{}N_{0}=(1+o(1))e^{-\nu}n\big{\}}$ (F.15)
@@
holds with $1-o(1)$ probability. In the following, for a generic event $\mathscr{E}$, an upper bound on $\mathbb{P}(\mathscr{E})$ in this part should be understood as a bound on $\mathbb{P}(\mathscr{E}\cap\mathscr{A}_{1})$, but we will only make this explicit in the concluding stage; see (F.52) below.
@@
We consider a fixed $(\ell,C,\zeta)$ and seek to establish the condition for (C2), i.e., (F.3) holding for some $\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}$ and some $\mathcal{J}^{\prime}\subset[p]\setminus s$ with $|\mathcal{J}^{\prime}|=\ell$. Note that $G_{\mathcal{J},\mathcal{J}^{\prime},1}$ in (F.3) is the number of positive tests in $\mathcal{N}_{0,\mathcal{J}^{\prime}}\cup\mathcal{M}_{\mathcal{J},\mathcal{J}^{\prime}}$, and thus we can further decompose $G_{\mathcal{J},\mathcal{J}^{\prime},1}$ into
$G_{\mathcal{J},\mathcal{J}^{\prime},1}=\widetilde{G}_{\mathcal{J}^{\prime},1}+U_{\mathcal{J},\mathcal{J}^{\prime},1},$ (F.16)
@@
where $\widetilde{G}_{\mathcal{J}^{\prime},1}$ is the number of positive tests in $\mathcal{N}_{0,\mathcal{J}^{\prime}}$ (with no dependence on $\mathcal{J}$, as reflected by the notation), and $U_{\mathcal{J},\mathcal{J}^{\prime},1}$ is the number of positive tests in $\mathcal{M}_{\mathcal{J},\mathcal{J}^{\prime}}$. Substituting (F.16) into (F.3), we can restate (C2) as follows:
$\widetilde{G}_{\mathcal{J}^{\prime},1}\geq\Big{(}\frac{1}{2}-\zeta\Big{)}\frac{Cn\nu e^{-\nu}\ell}{k}+\frac{1}{2}N_{0,\mathcal{J}^{\prime}}+\Big{(}\frac{1}{2}M_{\mathcal{J},\mathcal{J}^{\prime}}-U_{\mathcal{J},\mathcal{J}^{\prime},1}\Big{)}$ (F.17)
@@
$\max_{\begin{subarray}{c}\mathcal{J}^{\prime}\subset[p]\setminus s\\ |\mathcal{J}^{\prime}|=\ell\end{subarray}}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}-\frac{1}{2}N_{0,\mathcal{J}^{\prime}}\Big{)}\geq\Big{(}\frac{1}{2}-\zeta\Big{)}\frac{Cn\nu e^{-\nu}\ell}{k}-\max_{\begin{subarray}{c}\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}\\ \mathcal{J}^{\prime}\subset[p]\setminus s,|\mathcal{J}^{\prime}|=\ell\end{subarray}}M_{\mathcal{J},\mathcal{J}^{\prime}},$ (F.18)
@@
$\max_{\mathcal{J},\mathcal{J}^{\prime}}M_{\mathcal{J},\mathcal{J}^{\prime}}\leq\max_{\mathcal{J},\mathcal{J}^{\prime}}\widetilde{M}_{\mathcal{J},\mathcal{J}^{\prime}},$ (F.19)
where $(\mathcal{J},\mathcal{J}^{\prime})$ are implicitly subject to the same constraints as the last term in (F.18), and similarly in the developments below. We consider fixed $\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}$ and $\mathcal{J}^{\prime}\subset[p]\setminus s$ with $|\mathcal{J}^{\prime}|=\ell$. Since $\mathcal{M}_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k}$ and the test placements are uniform in $[n]$, a placement from items in $\mathcal{J}^{\prime}$ falls in $\mathcal{M}_{\mathcal{J}}$ with probability $\frac{C\nu e^{-\nu}\ell}{k}$, thus the randomness of $\mathbf{T}_{\mathcal{J}^{\prime}}$ gives $\widetilde{M}_{\mathcal{J},\mathcal{J}^{\prime}}\sim\text{Bin}(\frac{\ell\nu n}{k},\frac{C\nu e^{-\nu}\ell}{k})$. Then we can proceed analogously to the calculations in (D.28)–(D.31). In particular, our assumption $C\leq e^{\nu}$ is equivalent to $Ce^{-\nu}\leq 1$, and thus for given $(\mathcal{J},\mathcal{J}^{\prime})$ we have
$\mathbb{P}\Big{(}\widetilde{M}_{\mathcal{J},\mathcal{J}^{\prime}}\geq\frac{\ell n}{k(\log\frac{k}{\ell})^{1/2}}\Big{)}$ $\leq\mathbb{P}\Big{(}\text{Bin}\big{(}\frac{\nu\ell n}{k},\frac{\nu\ell}{k}\big{)}\geq\frac{\ell n}{k(\log\frac{k}{\ell})^{1/2}}\Big{)}$ (F.20)
$\leq\exp\Big{(}-\Omega\Big{(}\sqrt{\log\frac{k}{\ell}}\Big{)}\cdot\ell\log k\Big{)},$ (F.21)
where (F.21) follows from the same argument as the Bernoulli design (see (D.30)), with only the constant factors changing. Moreover, by repeating the union bound over $(\mathcal{J},\mathcal{J}^{\prime})$ as in (D.31), we obtain
$\mathbb{P}\Big{(}\max_{\mathcal{J},\mathcal{J}^{\prime}}\widetilde{M}_{\mathcal{J},\mathcal{J}^{\prime}}\geq\frac{\ell n}{k(\log\frac{k}{\ell})^{1/2}}\Big{)}\leq k^{-10\ell}$ (F.22)
@@
$\max_{\mathcal{J},\mathcal{J}^{\prime}}M_{\mathcal{J},\mathcal{J}^{\prime}}\leq\frac{\ell n}{k(\log\frac{k}{\ell})^{1/2}}=o\Big{(}\frac{\ell n}{k}\Big{)}.$ (F.23)
On this high-probability event, the necessary condition for (C2) given in (F.18) can be written as
$\max_{\begin{subarray}{c}\mathcal{J}^{\prime}\subset[p]\setminus s\\ |\mathcal{J}^{\prime}|=\ell\end{subarray}}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}-\frac{1}{2}N_{0,\mathcal{J}^{\prime}}\Big{)}\geq\Big{(}\frac{1}{2}-\zeta-o(1)\Big{)}\frac{Cn\nu e^{-\nu}\ell}{k}.$ (F.24)
Thus, to bound the probability of (C2), it suffices to bound the probability of its necessary condition (F.24). We proceed to study (F.24) by first studying a fixed $\mathcal{J}^{\prime}$ and then applying a union bound to account for the maximum.
@@
$\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}-\frac{1}{2}N_{0,\mathcal{J}^{\prime}}\geq\big{(}\frac{1}{2}-\zeta-o(1)\big{)}\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}\leq\mathbb{P}\Big{(}N_{0,\mathcal{J}^{\prime}}\leq\frac{d_{0}\nu e^{-\nu}\ell n}{k}\Big{)}$ (F.26)
@@
$+\sum_{\begin{subarray}{c}d_{0}<d\leq e^{\nu}\\ \frac{d\nu e^{-\nu}\ell n}{k}\in\mathbb{Z}\end{subarray}}\mathbb{P}\Big{(}N_{0,\mathcal{J^{\prime}}}=\frac{d\nu e^{-\nu}\ell n}{k}\Big{)}\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}\geq\frac{d+C(1-2\zeta)}{2}\frac{\nu e^{-\nu}\ell n}{k}\Big{|}N_{0,\mathcal{J}^{\prime}}=\frac{d\nu e^{-\nu}\ell n}{k}\Big{)}.$ (F.27)
@@
Case 1: $d_{0}\in[0,1]$ (i.e., $0\leq C(2\zeta-1)\leq 1-2\rho$). We will separately bound the probability terms in the right hand side of (F.26) and in (F.27).
@@
Bounding the term in (F.26): We first bound the term in the right-hand side of (F.26). By definition, ${N}_{0,\mathcal{J}^{\prime}}$ is the number of tests in $\mathcal{N}_{0}$ that contain some item from $\mathcal{J}^{\prime}$, thus on the event of $N_{0}=(1+o(1))ne^{-\nu}$ (as explained in the text below (F.15)), Lemma G.4 gives
$\mathbb{P}\Big{(}N_{0,\mathcal{J}^{\prime}}=\frac{d\nu e^{-\nu}\ell n}{k}\Big{)}$ $\leq\exp\Big{(}-\frac{\ell\nu n}{k}\big{[}D\big{(}de^{-\nu}\|e^{-\nu}+o(1)\big{)}+o(1)\big{]}\Big{)}$ (F.28)
$\leq\exp\Big{(}-\frac{\ell\nu n}{k}\big{[}D\big{(}d_{0}e^{-\nu}\|e^{-\nu}\big{)}+o(1)\big{]}\Big{)}$ (F.29)
@@
for any $d\leq d_{0}$ such that $\frac{d\nu e^{-\nu}\ell n}{k}\in\mathbb{Z}$, where (F.29) follows from $d_{0}\leq 1$ and the monotonicity of $D(de^{-\nu}\|e^{-\nu})$ with respect to $d$. Therefore, the term in the right-hand side of (F.26) can be bounded as
$\mathbb{P}\Big{(}N_{0,\mathcal{J}^{\prime}}\leq\frac{d_{0}\nu e^{-\nu}\ell n}{k}\Big{)}$ $=\sum_{0\leq d\leq d_{0}:\frac{d\nu e^{-\nu}\ell n}{k}\in\mathbb{Z}}\mathbb{P}\Big{(}N_{0,\mathcal{J}^{\prime}}=\frac{d\nu e^{-\nu}\ell n}{k}\Big{)}$ (F.30)
$\leq O\big{(}\frac{\ell n}{k}\big{)}\exp\Big{(}-\frac{\ell\nu n}{k}\big{[}D\big{(}d_{0}e^{-\nu}\|e^{-\nu}\big{)}+o(1)\big{]}\Big{)}$ (F.31)
$\leq\exp\Big{(}-\frac{\ell\nu n}{k}\big{[}D\big{(}d_{0}e^{-\nu}\|e^{-\nu}\big{)}+o(1)\big{]}\Big{)},$ (F.32)
@@
where (F.30) follows from $d_{0}=O(1)$ (see (F.25)) and hence there are at most $O(\frac{\ell n}{k})$ summands, and in (F.32) the leading factor $O(\frac{\ell n}{k})$ is absorbed into the $o(1)$ term in the exponent.
Bounding the term in (F.27): Analogously to (F.28), on the event of $N_{0}=(1+o(1))ne^{-\nu}$, Lemma G.4 gives
$\mathbb{P}\Big{(}N_{0,\mathcal{J}^{\prime}}=\frac{d\nu e^{-\nu}\ell n}{k}\Big{)}\leq\exp\Big{(}-\frac{\ell\nu n}{k}\cdot\big{[}D\big{(}de^{-\nu}\|e^{-\nu}\big{)}+o(1)\big{]}\Big{)}.$ (F.33)
@@
Then, conditioned on $N_{0,\mathcal{J}^{\prime}}=\frac{d\nu e^{-\nu}\ell n}{k}$ we have $\widetilde{G}_{\mathcal{J}^{\prime},1}\sim\text{Bin}(\frac{d\nu e^{-\nu}\ell n}{k},\rho)$ via the randomness of $\mathbf{Z}$, and thus the condition $d>d_{0}$ (which implies $\frac{d+C(1-2\zeta)}{2}>d\rho$) allows us to apply the Chernoff bound (see (A.1) in Lemma A.1) to obtain
$\mathbb{P}\Big{(}\text{Bin}\Big{(}\frac{d\nu e^{-\nu}\ell n}{k},\rho\Big{)}\geq\frac{d+C(1-2\zeta)}{2}\frac{\nu e^{-\nu}\ell n}{k}\Big{)}\leq\exp\left(-\frac{d\nu e^{-\nu}\ell n}{k}\cdot D\Big{(}\frac{1}{2}+\frac{C(1-2\zeta)}{2d}\big{\|}\rho\Big{)}\right).$ (F.34)